OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4,2) with a slope of 1/2. A.y=1/2 B.y=1/2x+4

OpenStudy (texaschic101):

y = mx + b 2 = 1/2(4) + b 2 = 2 + b 4 = b y = 1/2x + 4 < -- answer